8. For his business, Gil has determined that the time it takes to finish a job varies inversely with the number of workers. This can be expressed as the equation below where T = time, k is a constant, and w = number of workers.

Gil's records show that 18 workers can finish a job in 6 days.
How many days will it take 12 workers to do the same job?



A. 4
B. 9
C. 12
D. 36
1
Expert's answer
2012-11-07T08:38:58-0500
Let x is a number of days that 12 workers need tofinish this job. 12 workers by x days can do the same job as 18 workers by 6
days. the we can write the equation:
12*x=18*6
x=18*6/12=9
Answer: B. 9